# AI-Based Beam-Level and Cell-Level Mobility Management for High Speed Railway Communications

### v1

@@ III-A Use Cases in HSR Communications
Cell-level mobility management deals primarily with cell handovers, initiating the transition of a user’s session to a new cell when the signal strength from the serving cell falls below a predefined threshold. Given the rapid movement in HSR communications, cell handovers occur frequently, and failures in these handovers would unavoidable compromise the user experience. As specified in 3GPP, the cell handover procedure relies on concrete criteria, such as event A3, which will be triggered when the difference between the layer 3 RSRP (L3-RSRP)222L3-RSRP represents the cell-level RSRP, obtained by averaging or maximizing the RSRP of all beam pairs in each cell. measurement of the target cell and the L3-RSRP measurement of the source cell exceeds a predefined offset. If the above criteria are met during the time to trigger (TTT) duration, the UE reports the measurements and the source cell initiates the handover preparation phase. Afterwards, the cell handover procedure initiates the execution after the UE receives the handover command message, whose task is to provide radio resource control (RRC) reconfiguration message to the target cell. The specific cell handover process is shown in Fig. 4. During state 1, the UE must sweep across all beams to acquire the L3-RSRP of adjacent cells for the TTT duration, leading to signal quality deterioration in the service cell, which are even more pronounced in HSR scenarios.
To maintain a stable communication quality for HSR communications, the most common use case for AI-based cell-level mobility management is to predict the RSRP of all adjacent cells in future TTT instances. The AI-based cell handover process is illustrated in Fig. 4, where option 1 and option 2 represent different deployment methods of AI models. In this case, one option for the source cell is to immediately execute a cell handover based on the AI model’s predicted target cell. Alternatively, the source cell can monitor the signal quality of both the target and source cells over a period of time and initiate the handover if some handover conditions are still met. This approach can effectively reduce the ping-pong handovers and enhance the communication stability, while it requires more handover time for signal detection. In HSR scenarios, where trains move in straight lines and the probability of ping-pong handovers is low, the delay introduced by extended handover time, resulting in significant communication quality degradation, becomes intolerable. Therefore, the first approach is better suited for cell-level mobility management in HSR communications.
III-B Inputs, Outputs, and KPIs in HSR Communications
For the HSR scenario, we explore various model inputs and outputs tailored to its specific requirements. One input option for AI models is the L3-RSRP of each cell at historical instances, with the output being either the L3-RSRP of each cell or the L3-RSRP of the top cells at future TTT instances. The dimension of L3-RSRP information is lower than that of L1-RSRP. However, acquiring L3-RSRP, which requires averaging data over multiple time slots, introduces extra latency which is not desired in HSR communications. Another viable approach involves using layer 1 RSRP (L1-RSRP)333L1-RSRP represents the beam-level RSRP, which is the RSRP of each beam pair of each cell. as both the input and the output of the AI models. The AI model output is then filtered to derive the L3-RSRP, which is used to execute cell handovers. For this option, the output L1-RSRP of AI model can be applied to both beam-level and cell-level mobility management. Nevertheless, L1-RSRP encompasses extensive information that exceeds the requirements for handover, while the correlation between the input and output L1-RSRP is intricate, thereby amplifying the learning challenge for the AI models. Moreover, leveraging L1-RSRP to forecast L3-RSRP via AI can be an advantageous strategy. This method reaps the benefits of the above two approaches, making use of the abundant data in L1-RSRP to enhance the AI model’s prediction accuracy, while also streamlining the mapping between the model’s input and output.
For the evaluation of AI-based cell-level mobility management, we should not only focus on the prediction accuracy of RSRP, such as the normalized mean squared error (NMSE) or RSRP difference between the actual RSRP and the predicted RSRP, but we also need to pay more attention to the following system-level KPIs which are important for HSR communications.
RLF rates: Radio link failure (RLF) occurs when the signal-to-interference-plus-noise ratio (SINR) stays consistently small for an extended duration. In the HSR scenarios, the ultra-high moving speed necessitates timely handovers to mitigate SINR attenuation and reduce RLF rates. The RLF rates can be used to evaluate the timeliness of the cell handover during train operation.
Handover failure (HOF) rates: Handover failure usually happens due to too late handover, too early handover and handover to a wrong cell. The HOF rates is calculated as the total number of HOFs divided by the sum of the total number of HOFs and the total number of successful handover. The HOF rates is one indicator on the reliability of the HSR communication system.
Data interruption: The time interval between the transmission of the last data packet from the source BS and the reception of the first data packet from the target BS is referred to as data interruption. The low data interruption guarantees the timeliness of signal quality for high-speed trains.
Table I summarizes the enhancement schemes for AI-based mobility management.
| Ref | Category | Network | Input | Output | Description |
| [9] | Beam-level | FC | The sub-6G channel vector | The optimal beam and the path state like LoS | The model utilizes the low-frequency sub-6G channel to learn the mmWave channel information. |
| [10] | FC | Location information, mobility vectors, SNR and current beam | The link connectivity within the next beam tracking phase | Using an AI model to predict whether there will be a connection loss before the next beam tracking procedure. | |
| [11] | CNN | LIDAR measurements, the coordinates of the BS and UE | The optimal beam and the path state (LoS or NLoS) | Generating a histogram based on the model’s input and feeding it into a CNN to learn the channel features. | |
| [7] | CNN+FC | RSRP of partial narrow beams or all wide beams | The direction of the transceiver beams | The model utilizes supervised and unsupervised learning based on CNN and FC. | |
| [8] | CNN+FC | The RSRP of all beams of each UE | The optimal beam index of each UE | The model adopts a hybrid beamforming architecture and CNN to achieve multi-user synchronous beam alignment. | |
| [2] | LSTM | The complete RSRP images | The optimal beam index | The multi-layer LSTM network is utilized for temporal beam prediction, followed by FC for feature extraction. | |
| [6] | LSTM | Channel state information (CSI) and the context vector | CSI and beam quality | The LSTM-based prediction model is embedded into the extended Kalman filter (EKF) algorithm. | |
| [5] | CNN+LSTM | The RSRP images of all beam pairs | The optimal beam index of transceiver | The CNN is utilized for feature extraction, followed by LSTM for temporal beam prediction. | |
| [4] | Conv-LSTM | The RSRP image of all wide beams | The RSRP image of all narrow beams | The model use sub-pixel convolution and convolutional LSTM realizing image reconstruction. | |
| Our | CS+AI | Channel information of compressed multi-beam | RSRP of all beams at the current instance | Learning the optimal measurement matrix and conv-LSTM network to reduce the cost of beam sweeping. | |
| [3] | Cell-level | CNN+LSTM | SNR of all beams and all cells | The same as the input | The model utilizes CNN to reduce the input dimension and LSTM network to predict beam quality of all cells. |
| [15] | CNN/LSTM | RSRP images of all beams of all cells | The best BS index, and path state (LoS or NLoS) | Using two AI models based on CNN and LSTM predicts the best BS index and path state in future instances. | |
| Our | CNN/LSTM | Compressed RSRP images | RSRP of all beams of all cells | This article uses two AI models to predict the RSRP in future instances based on compressed RSRP images. |
III-C AI Enhancement for HSR Communications
@@
The traditional cell handover mechanism including 3 states is described in Fig. 4. In state 1, the UE measures the beam quality and sends the measurement report to the source BS. In this state, the UE remains connected to the source cell. When the mobility of the UE is low or the UE is moving within a macro cell, the above handover mechanism can sustain good performance. However, when the mobility of UE is high or the UE is moving across micro cells of high density, the traditional handover mechanism would result in unintended events, e.g., HOF, RLF, ping-pong phenomenon, throughput loss and too early/late handover. In HSR communications, the trains move at high speed, resulting in a sharp decline of the communication quality of the link between the UE and the source BS during state 1. Leveraging its robust environmental awareness and learning abilities, AI/ML technology can address the aforementioned issues.
@@
One approach for AI-based cell-level mobility management is to exploit the beam-level outputs of different cells. In particular, AI models are deployed at the network side, and each cell employs a unique AI model to predict the future RSRP of this cell. The shortcomings of this cell-level mobility management approach lie in the high deployment and management overhead of AI models. Another approach is to use an AI model to predict the future RSRP of multiple adjacent cells of the same type based on RSRP measurement information. In this context, [15] utilized a multi-layer LSTM architecture to predict the optimal BS index as the candidate for handover. Moreover, [3] developed a CNN-based auto-encoder to reduce the data dimension, and then a LSTM network was used to conduct link quality prediction for each cell. In literature, the input into these AI models is the RSRP of beam pairs of all cells, leading to significant measurement overhead and computational complexity.
@@
To this end, we consider two approaches to reduce the measurement overhead for cell-level mobility management. For the first approach, all beams from partially selected cells are used to predict the L3-RSRP for all cells. For the second approach, partially selected beams of each cell are used to predict the L3-RSRP for all cells. We utilize the partial L1-RSRP from the past six time instances to predict the L3-RSRP in the next four time instances. We employ CNN-based and LSTM-based models as shown in Fig. 5. The LSTM-based model comprises four LSTM layers and each layer has 168 hidden units. The features extracted by the LSTM network undergo further processing via an FC layer and a sigmoid function for feature mapping. To enhance the performance of the LSTM network, we adopt four identical networks to predict L3-RSRP for four future instances separately. For the CNN-based model, a set of regular 2D convolution layers with 12, 24 and 32 kernels are employed before feeding it into two FC layers. To avoid overfitting during the training process, max pooling layers are incorporated after each CNN layer.
@@
The channel used in our experiments is generated by QuaDRiGa, where we consider the UMa scenario and center frequency 30 GHz. We deploy 7 BSs, and each BS is equipped with three half-wavelength spaced uniform plane arrays (UPAs) with sector orientation of 30, 150 and -90 degrees. We generate four datasets corresponding to various UE speeds of 60 km/h, 120 km/h, 350 km/h, and 500 km/h. Each dataset comprises 50,000 samples, which are split into training, validation, and test sets in a 3:1:1 ratio. For the proposed two approaches with equidistant sampling, L1-RSRP of 50% beams in each cell or all beams in 50% cells are selected as input to the AI model. The batch size is 256 and the maximum epoch is 800. The learning rate dynamically changes from to , utilizing the adaptive moment estimation optimizer. The RSRP difference is one evaluation metric of prediction accuracy for cell-level mobility management, which reflects the disparity between the predicted L3-RSRP and the actual L3-RSRP of all cells. In addition, we also consider the RLF rates as the system-level evaluation metric to illustrate the overall handover performance of the communication system.
The RSRP difference between the AI predictions and the actual measurements is shown in Fig. 6. The experimental results show that both CNN-based and LSTM-based models show excellent cell-level temporal prediction performance. Table II shows the significantly reduced RLF rates in AI-assisted cell handover processes. Remarkably, the proposed two approaches to reduce the measurement overhead for cell-level beam management, labeled as Part_Cell and Part_Beam, achieve comparable RLF performance with the traditional approach using all beam measurements of all cells, labeled as All_Beam_Cell, while we can save 50% beam measurement overhead for HSR communications. Furthermore, as shown in Table II, the flops and total parameters of the CNN-based model are much lower than those of the LSTM-based model.
@@ IV-B Multi-Criteria Joint Decision-Making Mechanism for Mobility Management
In HSR environments, the rapid movement necessitates more adaptable mobility management strategies to effectively handle frequent environmental changes. Relying solely on RSRP variations is insufficient for ensuring communication stability during handover decisions. Therefore, a multi-criteria decision-making mechanism is imperative. This mechanism could integrate factors, such as signal interference levels, network load conditions, terminal velocity, and direction of movement, along with RSRP variations as inputs for AI models to obtain better handover decisions. This would greatly improve the accuracy and reliability, ensuring seamless communication for high-speed trains.
@@ IV-C Model Deployment and Compression
In practical scenarios, the AI-assisted mobility management methods require high energy dissipation due to the high computing complexity, so it is more meaningful to deploy these AI models on the network side. In this case, the UE reports measurement data to the source BS, which feeds the measurements into the AI/ML models for mobility management. Moreover, the transmission entails extra latency while the unavoidable transmission errors, occurring in reporting these measurements, affect the performance of AI models. Then, it would be beneficial to deploy AI models on the UE side. We can consider using CS or training lightweight AI models to reduce the computational complexity and hardware requirements for UEs.
@@ IV-D Model Generalization Design
AI/ML models designed for specific scenarios would usually have poor generalization performance. Altering the settings in communication scenarios, e.g., BS height, channel model, cell size and UE speed, might lead to a significant deterioration of the model performance. Therefore, one challenge for AI-based mobility management for HSR communications is the design of AI models with good scenario generalization properties. Generally, there exists a trade-off between model generalization and model performance. Improving the generalization of the model, e.g., training the model with a mixed datasets from multiple scenarios, will introduce difficulties for the training of AI/ML models, as more complex and diverse knowledge are embedded in the data. As a result, the balance between model generalization and prediction accuracy according to the actual application scenario needs to be thoroughly examined in future discussions.
@@ V Conclusion
In this article, we investigated the fundamental issue of AI-based beam-level and cell-level mobility management for HSR communications, including the use cases, inputs, outputs and KPIs of AI models. For beam-level mobility management in HSR communications, a hybrid approach, combining AI and CS, demonstrated superior performance compared to conventional AI-based methods with the same measurement overhead. For cell-level mobility management in HSR communications, we unveiled the significant gain of AI-based methods for cell handover, while the measurement overhead can be reduced by using partial beam measurements, without compromise of the RLF rates. At last, the opportunities and challenges on AI-based mobility management are briefly discussed for further studies.

### v2

@@ III-A AI-Based Cell-Level Mobility Management in HSR Communications
Cell-level mobility management deals primarily with cell handovers, initiating the transition of a user’s session to a new cell when the signal strength from the serving cell falls below a predefined threshold. Given the rapid movement in HSR communications, cell handovers occur frequently, and failures in these handovers would unavoidably compromise the user experience. Handover decision is performed by the source cell based on UE measurement. UE measurement should be met by certain criteria, for example, event A3 which will be triggered when the difference between the layer 3 RSRP (L3-RSRP)444L3-RSRP represents the cell-level RSRP, obtained by averaging or maximizing the RSRP of all beam pairs in each cell. measurement of the target cell and the L3-RSRP measurement of the source cell exceeds a predefined offset. If the above criteria are met during the time to trigger (TTT) duration, the UE reports the measurements and the source cell initiates the handover preparation phase. Afterwards, the cell handover procedure initiates the execution after the UE receives the handover command message, whose task is to provide RRC reconfiguration message to the target cell. The specific cell handover process is shown in Fig. 5.
@@
In state 1, the UE measures the beam quality and sends the measurement report to the source BS. In this state, the UE remains connected to the source cell. When the mobility of the UE is low or the UE is moving within a macro cell, the traditional handover mechanism can sustain good performance. However, when the mobility of UE is high or the UE is moving across micro cells of high density, the traditional handover mechanism would result in unintended events, e.g., handover failure, radio link failure (RLF), ping-pong phenomenon, throughput loss. In HSR communications, the trains move at high speed, resulting in a sharp decline of the communication quality of the link between the UE and the source BS during state 1. Leveraging its robust environmental awareness and learning abilities, AI/ML technology can address the aforementioned issues.
@@
One approach for AI-based cell-level mobility management is to exploit the beam-level outputs of different cells. In particular, AI models are deployed at the network side, and each cell employs a unique AI model to predict the future RSRP of this cell. The shortcomings of this cell-level mobility management approach lie in the high deployment and management overhead of AI models. Another approach which is more suitable for HSR scenarios is to use an AI model to predict the future RSRP of multiple adjacent cells or the future optimal BS index based on historical RSRP measurement information. In this context, [15] utilized a multi-layer LSTM architecture to predict the optimal BS index as the candidate for handover. Moreover, [3] developed a CNN-based auto-encoder to reduce the data dimension, and then a LSTM network was used to conduct link quality prediction for each cell.
For the HSR scenario, we also explore various model inputs and outputs tailored to its specific requirements. One input option for AI models is the L3-RSRP of each cell, with the output being either the L3-RSRP of each cell or the L3-RSRP of the top cells. The dimension of L3-RSRP information is lower than that of layer 1 RSRP (L1-RSRP)555L1-RSRP represents the beam-level RSRP, which is the RSRP of each beam pair of each cell. However, acquiring L3-RSRP, which requires averaging data over multiple time slots, introduces extra latency which is not desired in HSR communications. Another viable approach involves using L1-RSRP as both the input to and the output of the AI models. The AI model output is then filtered to derive the L3-RSRP, which is used to execute cell handovers. For this option, the output L1-RSRP of the AI model can be applied to both beam-level and cell-level mobility management. Nevertheless, L1-RSRP encompasses extensive information that exceeds the requirements for handover, while the correlation between the input and output L1-RSRP is intricate, thereby amplifying the learning challenge for the AI models. Moreover, leveraging L1-RSRP to forecast L3-RSRP via AI can be an advantageous strategy. This method reaps the benefits of the above two approaches, making use of the abundant data in RSRP to enhance the AI model’s prediction accuracy, while also streamlining the mapping between the model’s input and output.
Ref | Category | Network | Input | Output | Description | KPIs |
[11] | Beam-level | FC | The sub-6G channel vector | The optimal beam and the path state like LoS | Utilizing the low-frequency sub-6G channel to predict. | Top-1 accuracy, spectral efficiency |
[13] | FC | Location information, mobility vectors, SNR and current beam | The link connectivity within the next beam tracking phase | Predicting whether there will be a connection loss before the next beam tracking procedure. | Throughput | |
[12] | CNN | LIDAR measurements, the coordinates of the BS and UE | The optimal beam and the path state: line-of-sight (LoS) or non line-of-sight (NLoS) | Generating a histogram based on the model’s input to learn the channel features. | Accuracy for LoS detection | |
[9] | CNN+FC | RSRP of partial narrow beams or all wide beams | The direction of the transceiver beams | Utilizing supervised and unsupervised learning based on CNN and FC. | Achievable rate | |
[10] | CNN+FC | The RSRP of all beams of each UE | The optimal beam index of each UE | Adopting a CNN to achieve multi-user synchronous alignment. | Spectral efficiency | |
[2] | LSTM | The complete RSRP images | The optimal beam index | The multi-layer LSTM network and FC are utilized. | RSRP difference,Top-K accuracy | |
[8] | LSTM | Channel state information (CSI) and the context vector | CSI and beam quality | The LSTM network and extended Kalman filter (EKF) are used. | Achievable rate | |
[7] | CNN+LSTM | The RSRP images of all beam pairs | The optimal beam index of transceiver | Utilizing CNN for feature extraction and LSTM for temporal beam prediction. | RMSE | |
[6] | Conv-LSTM | The RSRP image of all wide beams | The RSRP image of all narrow beams | Using sub-pixel convolution and Conv-LSTM to realize RSRP prediction. | MSE | |
Our | CS+AI | Channel information of compressed multi-beam | RSRP of all beams at the current instance | Learning the optimal measurement matrix and conv-LSTM network for prediction. | Top-1 accuracy | |
[3] | Cell-level | CNN+LSTM | SNR of all beams and all cells | The same as the input | Using CNN to reduce dimension and LSTM network to predict beam quality. | link(s) prediction error |
[15] | CNN/LSTM | RSRP images of all beams of all cells | The best BS index, and path state (LoS or NLoS) | Using two AI models based on CNN and LSTM predicts the best BS index. | Prediction accuracy of best BS | |
Our | CNN/LSTM | Compressed RSRP images | RSRP of all beams of all cells | Using two AI models to optimize the shortcomings of the handover mechanism. | RLF rates |
III-B AI-Based Proactive Handover Scheme
Existing AI/ML-assisted cell handover methods typically use the RSRP of all beams across cells to predict the future RSRP for each beam or to identify the optimal base station index for upcoming handovers [15, 3]. These methods primarily focus on the improvements in prediction accuracy and spectral efficiency provided by AI models, with limited attention to system-level metrics, such as RLF rates and PP rates. This gap may overlook the potential impact of model design on system reliability and service continuity. Additionally, existing algorithms do not sufficiently address the inherent limitations of the event A3-based handover mechanism. For instance, event A3-based handover is triggered and executed based on historical measurement reports, making it a passively reactive approach. The system initiates the handover only after the UE detects communication quality degradation, where the communication quality has already suffered irreversible loss.
To maintain a stable communication quality for HSR communications, we provide a novel AI-assisted cell handover scheme. We use an AI model to predict the L3-RSRP of all cells for future state 1 based on historical L1-RSRP to aid the source BS in handover decisions. Using this method, we predict the future communication quality variation trends of adjacent cells and proactively switch cells based on these predictions. This approach effectively mitigates the communication quality degradation and reduces handover time from the original states (state 1, state 2, state 3) to the existing states (state 2, state 3). The AI-based cell handover process is illustrated in Fig. 5, where option 1 and option 2 represent different deployment methods of AI models. In option 1, the AI-based prediction model is deployed on the UE side, with the prediction report uploaded to the source BS to assist in the handover decision. In option 2, the AI-based prediction model is deployed on the network side, where the network performs temporal prediction and makes the handover decision based on historical measurement reports. The primary difference between the two options lies in the model’s deployment mode, while the model’s inputs, outputs, and architecture remain the same.
@@ III-B AI-Based Proactive Handover Scheme
Based on the handover scheme we propose, we identify two approaches to reduce measurement overhead in cell-level mobility management. For the first approach, all beams from the partially selected cells are used to predict the L3-RSRP for all cells. For the second approach, the partially selected beams of each cell are used to predict the L3-RSRP for all cells. To be noticed, the model input is obtained through equal-interval sampling rather than the compressed multi-beam measurement method proposed in Section II to independently evaluate the performance gains of the AI-assisted handover scheme. We utilize the partial L1-RSRP from the past six time instances to predict the L3-RSRP in the next four time instances. We employ CNN-based and LSTM-based models for RSRP prediction. The LSTM-based model comprises four LSTM layers and two FC layers. To enhance the performance of the LSTM network, we adopt four identical networks to predict L3-RSRP for four future instances separately. For the CNN-based model, a set of regular 2D convolution layers with 12, 24 and 32 kernels are employed before feeding it into two FC layers. To avoid overfitting during the training process, max pooling layers are incorporated after each CNN layer.
@@
The channel used in our experiments is generated by QuaDRiGa, with the UMa scenario and a center frequency of 30 GHz, consistent with the beam-level experiment setup. We deploy 7 BSs, and each BS is equipped with three half-wavelength spaced uniform plane arrays (UPAs) with sector orientation of 30, 150 and -90 degrees and interference is generated by 21 adjacent cells, modeled as random interference. We generate four datasets corresponding to various UE speeds of 60 km/h, 120 km/h, 350 km/h, and 500 km/h. Each dataset comprises 50,000 samples, which are split into training, validation, and test sets in a 3:1:1 ratio. For the proposed two approaches with equidistant sampling, L1-RSRP of 50% beams in each cell or all beams in 50% cells are selected as input to the AI model. The batch size is 256 and the maximum epoch is 800. The learning rate dynamically changes from to , utilizing the adaptive moment estimation optimizer. The RSRP difference is one evaluation metric of prediction accuracy for cell-level mobility management, which reflects the disparity between the predicted L3-RSRP and the actual L3-RSRP of all cells. In addition, we also consider the RLF rates as the system-level evaluation metric to illustrate the overall handover performance of the communication system. The experimental results show that both CNN-based and LSTM-based models show excellent cell-level temporal prediction performance. Table II shows the significantly reduced RLF rates in AI-assisted cell handover processes. Remarkably, the proposed two approaches to reduce the measurement overhead for cell-level beam management, labeled as Part_Cell and Part_Beam, achieve comparable RLF performance with the traditional approach using all beam measurements of all cells, labeled as All_Beam_Cell, while we can save 50% beam measurement overhead for HSR communications. Furthermore, as shown in Table II, the flops and total parameters of the CNN-based model are much lower than those of the LSTM-based model, while achieving comparable performance.
@@ IV-B Multi-Criteria Joint Decision-Making Mechanism for Mobility Management
In HSR environments, the rapid movement necessitates more adaptable mobility management strategies to effectively handle frequent environmental changes. Relying solely on RSRP variations is insufficient for ensuring communication stability during handover decisions. Therefore, a multi-criterion decision-making mechanism becomes imperative. This mechanism could integrate factors, such as signal interference levels, network load conditions, terminal velocity, and direction of movement, along with RSRP variations as inputs to the AI models to obtain better handover decisions. This would greatly improve the accuracy and reliability, ensuring seamless communication for high-speed trains.
@@ IV-C Model Deployment and Compression
In practical scenarios, the AI-assisted mobility management methods require high energy dissipation due to the high computing complexity, so it is more meaningful to deploy these AI models on the network side. In this case, the UE reports measurement data to the source BS, which feeds the measurements into the AI/ML models for mobility management. Moreover, the transmission entails extra latency while the unavoidable transmission errors, occurring during the reporting of these measurements, affect the performance of AI models. Then, it would be beneficial to deploy AI models on the UE side. We can consider employing advanced AI models, such as lightweight AI models or transformer-based architectures, to enhance the computational efficiency and reduce the hardware requirements for the UEs.
@@ IV-D Model Generalization Design
AI/ML models designed for specific scenarios would usually have poor generalization performance. Altering the settings in communication scenarios, e.g., BS height, channel model, cell size and UE speed, might lead to a significant deterioration of the model performance. Therefore, one challenge for AI-based mobility management for HSR communications is the design of AI models with good scenario generalization properties. Continuous learning is particularly beneficial in dynamic environments, enabling models to adapt to shifting channel conditions and movement patterns in real time. However, enhancing the model generalization through continuous learning incurs significant deployment overhead. Balancing deployment overhead and generalization remains a critical research in HSR communication.
@@ V Conclusion
In this article, we investigated the fundamental issue of AI-based beam-level and cell-level mobility management for HSR communications of AI models. For beam-level mobility management in HSR communications, a hybrid approach, combining AI and CS, demonstrated superior performance compared to conventional AI-based methods with the same measurement overhead. For cell-level mobility management in HSR communications, we unveiled the significant gain of AI-based methods for cell handover, while the measurement overhead can be reduced by using partial beam measurements, without compromise of the RLF rates. At last, the opportunities and challenges on AI-based mobility management were briefly discussed for further studies.